8.) What is the chi square equation?

Select one:

a. Χ2 = Σ [(O – E)2/E]

b. Χ3 = Σ [(O – E)3/E]

c. Χ2 = Σ [(O – E)2*E]

d. Χ2 = Σ [(O – E)3/E]

9.) Calculate the degrees of freedom if 5 defined classes are being rested. Show your work.

10.) Determine the X2 using the following information. Show your work.

You hypothesize that the gene for rat coat color follows Mendelian inheritance. After you perform a cross between two monohybrids, you expect to get 75 black coats and 25 white coats in a population of 100. However, you observe 71 black coats and 29 white coats.

11.) Using your answer from #10, determine the p value. Is your hypothesis supported or rejected? How do you know?

24.)  

Suppose you perform a monohybrid cross of two pea plants. The plants are heterozygous at the locus for stem height. T is dominant to t. T confers tall plant height. When you count the progeny, you find that there are 202 tall plants and 98 short plants. Does this approximate the 3:1 phenotypic ratio caused by Mendelian inheritance?

Select one:

a. No, you would expect there to be more short plants.

b. Yes, there are approximately 3 tall plants for every short plant.

c. Yes, there are approximately 3 short plants for every tall plant.

d. No, the number of each phenotype is not in the exact ratio of 3:1.
Answer #1

8)

Chi square equation= £ { ( O- E) ^2/E}

Right option is A.

9) There are 5 classes. Here n value is 5

Degree of freedom = n-1

= 5-1 = 4

Please look in the image for solution of question 10 and 11)

0-E Expected bser 2-285 0.064 25 2 2.925 Thats whu OSi